# Universal Kerr-thermal dynamics of self-injection-locked microresonator dark pulses

### v1

@@
Shichang Li International Quantum Academy, Shenzhen 518048, China Shenzhen Institute for Quantum Science and Engineering, Southern University of Science and Technology, Shenzhen 518055, China    Kunpeng Yu International Quantum Academy, Shenzhen 518048, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China    Dmitry A. Chermoshentsev Russian Quantum Center, Moscow 143026, Russia    Wei Sun sunwei@iqasz.cn International Quantum Academy, Shenzhen 518048, China    Jinbao Long International Quantum Academy, Shenzhen 518048, China    Xiaoying Yan International Quantum Academy, Shenzhen 518048, China    Chen Shen International Quantum Academy, Shenzhen 518048, China Qaleido Photonics, Shenzhen 518048, China    Artem E. Shitikov Russian Quantum Center, Moscow 143026, Russia    Nikita Yu. Dmitriev Russian Quantum Center, Moscow 143026, Russia    Igor A. Bilenko Russian Quantum Center, Moscow 143026, Russia Faculty of Physics, Lomonosov Moscow State University, Moscow 119991, Russia    Junqiu Liu liujq@iqasz.cn International Quantum Academy, Shenzhen 518048, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China
@@
Microcombs, formed in optical microresonators driven by continuous-wave lasers, are miniaturized optical frequency combs with small size, weight and power consumption. Leveraging integrated photonics and laser self-injection locking (SIL), compact and robust microcombs can be constructed via hybrid integration of a semiconductor laser with a chip-based microresonator. While the current linear SIL theory has successfully addressed the linear coupling between the laser cavity and the external microresonator, it fails to manage the complicated nonlinear processes, especially regarding to dark-pulse microcomb formation. Here, we investigate – theoretically, numerically and experimentally – the Kerr-thermal dynamics of a semiconductor laser self-injection-locked to an integrated silicon nitride microresonator. We unveil intriguing yet universal dark-pulse formation and switching behaviour with discrete steps, and establish a theoretical model scrutinizing the synergy of laser-microresonator mutual coupling, Kerr nonlinearity, photo-thermal effect. Numerical simulation confirms the experimental result and identifies the origins. Exploiting this unique phenomenon, we showcase an application on low-noise photonic microwave generation with phase noise purified by 23.5 dB. Our study not only add critical insight of pulse formation in laser-microresonator hybrid systems, but also enables all-passive, photonic-chip-based microwave oscillators with high spectral purity.
@@
Optical frequency combs (OFCs) Udem et al. (2002); Cundiff and Ye (2003); Fortier and Baumann (2019); Diddams et al. (2020) constitute broadband frequency rulers with equidistant lines of continuous-wave (CW) components, and have revolutionized timing, spectroscopy, precision measurement, and testing fundamental physics. While conventional OFCs are constructed with solid-state or fiber mode-locked lasers, microcombs – harnessing enhanced Kerr nonlinearity in ultrahigh-$Q$ optical microresonators driven by CW pumps – have allowed miniaturized OFCs with small size, weight and power consumption Del’Haye et al. (2007); Kippenberg et al. (2018). With the emergence and quick maturing of ultralow-loss photonic integrated circuits Moss et al. (2013); Gaeta et al. (2019); Zhang et al. (2017); Liu et al. (2021), as well as heterogeneous and hybrid integration with semiconductor lasers Stern et al. (2018); Xiang et al. (2021), today microcombs can be built entirely on-chip and manufactured in large volume with low cost, catalyzing wide deployment outside laboratories and in space.
@@
Depending on the microresonator’s group velocity dispersion (GVD), there are two types of coherent microcombs. The bright dissipative soliton microcombs Herr et al. (2013); Yi et al. (2015); Brasch et al. (2016a); Liang et al. (2015); Joshi et al. (2016); He et al. (2019) require anomalous GVD, while the dark-pulse (also termed “platicon”) microcombs require normal GVD Xue et al. (2015); Lobanov et al. (2015); Huang et al. (2015); Parra-Rivas et al. (2016); Nazemosadat et al. (2021); Wang et al. (2022). Compared to solitons, platicons exhibit remarkably higher CW-to-pulse power conversion efficiency Xue et al. (2017); Jang et al. (2021), thus are advantageous for coherent optical communication Fülöp et al. (2018) and photonic microwave generation Sun et al. (2024a). To initiate platicons, laser self-injection locking (SIL) Kondratiev et al. (2017, 2023) offers the most robust and effective form, and permits seamless integration of the pump laser and the microresonator Voloshin et al. (2021); Jin et al. (2021); Lihachev et al. (2022). However, while the current linear SIL theory and model Kondratiev et al. (2017) have addressed the linear coupling between the laser cavity and the external microresonator, it fails to manage the complicated nonlinear processes in the laser gain media and in platicon formation.
@@
Figure 1: Principle and schematic of self-injection-locking, Kerr nonlinearity and photo-thermal effect in the laser-microresonator coupled system. a. Frequency-domain picture illustrating the SIL process with the Kerr-thermal effect. b. Frequency- and time-domain pictures describing the synergy of SIL, photo-thermal effect and Kerr nonlinearity, which yields narrowing of the pump laser’s linewidth, red-shift of the microresonator’s resonance grid, and formation of a platicon microcomb. c. Experimental setup showing the layout of a DFB laser chip edge-coupled to a Si3N4 microresonator chip. A home-made printed circuit board (PCB) is used to provide stable current to the laser and to stabilize the laser temperature. Output light from the Si3N4 microresonator is collected by a lensed fiber. d. Schematic of SIL. $E_{\text{L}}$: light amplitude in the laser cavity. $E^{-}$/$E^{+}$: light amplitude in the clockwise/counter-clockwise direction in the microresonator. $E_{\text{out}}$: light amplitude at the microresonator output. $\varphi$: feedback phase of $E^{-}$.
@@
Here, we investigate – theoretically, numerically and experimentally – the Kerr-thermal dynamics of platicon formation using a semiconductor laser self-injection-locked to an integrated silicon nitride (Si3N4) microresonator. We unveil an intriguing platicon switching behaviour with discrete steps, allowing operation of platicons in noise-quenched states and immunity to laser noise. Figure 1a illustrates the principle of our study. When a laser-cavity mode is tuned into a resonance mode of the Si3N4 microresonator and light is coupled into the microresonator, SIL occurs that locks the laser frequency to the resonance Kondratiev et al. (2017, 2023) and suppresses the laser linewidth. The linewidth suppression ratio is proportional to $Q_{\text{r}}/Q_{\text{d}}$, where $Q_{\text{r/d}}$ is the quality factor of the microresonator/laser cavity. Meanwhile, photo-thermal effect Carmon et al. (2004); Brasch et al. (2016b); Yi et al. (2016); Gao et al. (2022) induces a global frequency shift of the microresonator’s resonance grid. With sufficient intracavity power, Kerr nonlinearity induces four-wave mixing (FWM) Kippenberg et al. (2004) that translates photons to other resonances. In microresonators of normal GVD, the synergy of SIL, photo-thermal and Kerr effects ultimately yield platicon formation, as shown in Fig. 1b.
@@
Experimental result. We use a commercial DFB laser chip edge-coupled to a Si3N4 microresonator chip Ye et al. (2023); Sun et al. (2024a), as shown in Fig. 1c. The DFB laser operates at 1549 nm in the telecommunication C band. It has 1 nm wavelength tuning range and 159 mW output CW power with driving current up to 500 mA. The Si3N4 microresonator has 10.7 GHz free spectral range (FSR) and normal GVD. ($D_{2}<0$). The microresonator is critically coupled Pfeiffer et al. (2017), with intrinsic quality factor $Q_{0}=23\times 10^{6}$. Actual photographs and characterization data of the DFB laser chip and the Si3N4 microresonator chip are found in Supplementary Information Note 1.
@@
Figure 1d illustrates the schematic of SIL. Light $E_{\text{L}}$ in the laser cavity is emitted and coupled into the microresonator’s counter-clockwise direction $E^{+}$. Rayleigh scattering in the microresonator reflects a portion of light to the clockwise direction $E^{-}$ and to the laser cavity. To facilitate SIL, we optimize the edge coupling and the gap distance (thus the feedback phase $\phi$) between the two chips. In the SIL regime, we continuously tune the laser current, and monitor the output optical power and frequency from the Si3N4 chip. The output light is beaten against a reference laser, and the beat frequency is recorded. The forward (backward) tuning corresponds to increasing (decreasing) laser current, which decreases (increases) laser frequency.
@@
Figure 2(a, b) shows the experimental data. When SIL occurs, the output power experiences a sudden drop, and the output frequency is shifted. In the SIL regime, we observe decreasing (increasing) output power and frequency with backward (forward) tuning, due to the photo-thermal effect Carmon et al. (2004); Brasch et al. (2016b); Yi et al. (2016); Gao et al. (2022). For example, the decreasing output power corresponds to decreasing transmitted power and increasing intra-cavity power. The latter enhances the photo-thermal effect and causes increasing resonance red-shift (towards longer wavelength) Carmon et al. (2004). As the laser frequency is locked to the resonance, the red-shifted resonance drags the laser frequency, leading to decreasing output optical frequency.
@@
Besides, it is apparent that the output optical power and frequency exhibit discrete step features. This is contrary to the conventional, linear SIL model Kondratiev et al. (2017) showing continuous, nearly anchored tuning curves of power and frequency. Here, the appearance of these steps are attributed to the Kerr nonlinearity in the microresonator with sufficient intra-cavity power. Note that similar discontinuous curves have been observed and characterized in self-injection-locked solitons Voloshin et al. (2021), while here we observe and characterize these curves for the first time for self-injection-locked platicons. Figure 2c shows the zoom-in profile of the gray-shaded zoom in Fig. 2b. Steps in backward tuning with 320.4 to 321.3 mA laser current are marked with 1 to 4. Typical optical spectra within each step are shown in Fig. 3a, evidencing formation and switching of different platicon states. Distinct fringes are marked with arrows on the spectral envelopes. Numerical simulation of time-domain pulse shapes in Fig. 3c (discussed later) confirms that all these platicon states are “single platicon” comprising only one dark pulse in the microresonator.
@@
Moreover, identical Kerr-thermal platicon dynamics has also been observed in other two independent SIL setups, where different DFB lasers and 21.3-GHz-FSR Si3N4 microresonators are used. Details are found in Supplementary Information Note 1 and 2. These parallel experiments suggest that, our observation of the Kerr-thermal dynamics and discrete steps is universal and independent of the particular lasers or microresonators.
@@
Figure 2: Experimental result in comparison with numerical simulation. a. Experimentally measured output laser power with forward (green) or backward (yellow) tuning. b. Experimentally measured beat frequency between the output laser and a frequency-fixed reference laser, with forward (blue) or backward (red) tuning. c. Zoom-in profiles of the gray-shaded region in b with forward (blue) or backward (red) tuning. Discrete steps with backward tuning are numbered with 1 to 4. d, e. Numerical simulation results corresponding to experimental data in a, b. Horizontal axes are the frequency detuning $\xi$ (in the unit of $\kappa/2$) between the free-running laser frequency and the cold resonance frequency. Effective (Eff.) detuning is the frequency detuning (in the unit of $\kappa/2$) between the locked laser frequency and the cold resonance frequency. f. Zoom-in profile of the gray-shaded zoom in d, e. Gray curves outlines the full detuning range of platicon steps.
@@
Theory and numerical simulation. To better understand our experimental result, we establish a comprehensive theoretical model harmonizing SIL, Kerr nonlinearity, and photo-thermal effect. The equations describing our laser-microresonator nonlinear system are written as
$\partial_{\tau}g_{N}$ $=J_{N}-\tilde{\kappa}_{N}(g_{N}+g_{0})|a_{\text{L}}|^{2}$ (1)
@@
$\partial_{\tau}a_{\text{L}}$ $=\frac{1}{2}\left(1+i\alpha_{g}\right)g_{N}a_{\text{L}}+i\mathrm{e}^{i(\alpha_{\text{L}}\tau+\varphi)}\tilde{\kappa}_{\text{L}}b_{0}$ (2)
@@
$\partial_{\tau}a_{\mu}$ $=-\left(1+i\sum_{n=2}^{\infty}\frac{d_{n}}{n!}\mu^{n}-iT\right)a_{\mu}+i\left(\mathcal{F}[|a|^{2}a]_{\mu}+2a_{\mu}\sum_{\mu^{\prime}}\left|b_{\mu^{\prime}}\right|^{2}\right)+i\tilde{\beta}b_{\mu}+i\mathrm{e}^{i(-\alpha_{\text{L}}\tau+\varphi)}\bar{\kappa}_{\text{R}}a_{\text{L}}\delta_{0\mu}$ (3)
@@
$\partial_{\tau}T$ $=\left(k_{T}\sum_{\mu}\left(\left|a_{\mu}\right|^{2}+\left|b_{\mu}\right|^{2}\right)-T\right)/t_{\text{h}}.$ (5)
@@
Derivation of Eq. (1–5) is found in Supplementary Information Note 3. In short, Eq. (1) describes the time evolution of the electron carrier inside the laser cavity, with $\tau=\kappa t/2$, $g_{N}=2G_{n}(N-N_{0})/\kappa$, $J_{N}=2I/(\kappa eV_{\text{L}})$, $\tilde{\kappa}_{N}=3G_{n}/(gV_{\text{L}})$, $g_{0}=2\tilde{G}_{0}/\kappa$, and $a_{\text{L}}=\sqrt{g\epsilon_{0}n_{\text{L}}^{2}V_{\text{L}}/(\hbar\omega_{\text{L}}\kappa)}E_{\text{L}}$. Here, $\kappa$ is the loaded linewidth of the Si3N4 microresonator, $t$ is the real time, $G_{n}$ is the gain coefficient of the active area in the laser cavity, $N$ is the electron carrier density, $N_{0}$ is the electron carrier density of the free-running laser with a fixed driving current, $I$ is the driving current on the laser diode, $e$ is the electron charge, $V_{\text{L}}$ is the effective volume of the laser diode’s active area, $\tilde{G}_{0}$ is the net gain at $N_{0}$ of the gain media in the laser, $\epsilon_{0}$ is the vacuum permittivity, $n_{\text{L}}$ is the refractive index of the active area, $\omega_{\text{L}}$ is the laser cavity resonance frequency, and $E_{\text{L}}$ is the laser amplitude in the laser cavity. The nonlinear coefficient of the Si3N4 microresonator is $g=\hbar\omega_{0}^{2}cn_{2}/(n_{0}^{2}V_{\text{eff}})$, where $\hbar$ is the reduced Plank constant, $\omega_{0}$ is the microresonator resonance frequency, $c$ is the speed of light in vacuum, $n_{2}$ is the nonlinear refractive index of Si3N4, $n_{0}$ is the refractive index of Si3N4, and $V_{\text{eff}}$ is the effective volume of the Si3N4 microresonator.
@@
Equation (2) describes the laser dynamics in the laser cavity, with $\alpha_{\text{L}}=2\delta\omega_{\text{L}}/\kappa$, $\varphi=\omega_{0}t_{\text{s}}$, $\tilde{\kappa}_{\text{L}}=2\sqrt{\kappa_{\text{ex}}\tau_{\text{R}}T_{\text{L}}T_{\text{c}}n_{\text{L}}V_{\text{L}}}/(\kappa\tau_{\text{L}}\sqrt{n_{0}V_{\text{eff}}})$. Here, $\alpha_{g}$ is the linewidth enhancement factor, $\delta\omega_{\text{L}}$ is the frequency difference between the cold-cavity resonance mode and the microresonator resonance mode, $t_{\text{s}}$ is the time delay between the laser cavity and the microresonator, $\kappa_{\text{ex}}$ is the external coupling rate of the microresonator, $\tau_{\text{R}}$ is the light round-trip time in the microresonator, $T_{\text{L}}$ is the light transmission rate at the laser cavity’s output surface, $T_{\text{c}}$ is the light coupling rate into the laser cavity from outside, $\tau_{\text{L}}$ is the light round-trip time in the laser cavity, and $b_{0}$ is the back-scattered light in the microresonator at the $0^{\text{th}}$ resonance mode.
@@
Equations (3, 4) describe the dynamics of the forward-propagating light $a_{\mu}$ and the back-scattered light $b_{\mu}$ at the $\mu^{\text{th}}$ resonance mode, with $a_{\mu}(b_{\mu})=\sqrt{g\epsilon_{0}n_{0}^{2}V_{\text{eff}}/(\hbar\omega_{0}\kappa)}E_{\mu}^{+}(E_{\mu}^{-})$, $d_{n}=2D_{n}/\kappa$, $\tilde{\beta}=2\beta/\kappa$, and $\tilde{\kappa}_{\text{R}}=2\sqrt{\kappa_{\text{ex}}T_{\text{L}}T_{\text{c}}n_{0}V_{\text{eff}}}/(\kappa\sqrt{\tau_{\text{R}}n_{\text{L}}V_{\text{L}}})$. Here, $E_{\mu}^{+}(E_{\mu}^{-})$ is the laser amplitude of the forward-propagating (back-scattered) light at the $\mu^{\text{th}}$ resonance mode, $D_{n}$ is the $n^{\text{th}}$-order dispersion of the microresonator, $\beta$ is the back-scattering ratio between the back-scattered light and the forward-propagating light, and $\mathcal{F}[|a|^{2}a]$ ($\mathcal{F}[|b|^{2}b])$ represents the Kerr interaction.
@@
Equation (5) describes the evolution of the normalized frequency shift $T$ induced by the temperature in the microresonator, with $k_{T}=K_{T}/g$, and $t_{\text{h}}=\kappa\tau_{\text{r}}/2$. Here, $K_{T}$ is the thermal-induced resonance shift coefficient, and $\tau_{\text{r}}$ is the thermal relaxation time.
@@
Figure 3: Optical spectra of different platicon states and the resulted noise-quenching effect. a, b. Experimentally measured a and simulated b optical spectra of the platicon states labelled in Fig. 2c, f. They agree not only on the spectral envelopes but also on the number of fringes (marked with arrows). c. Simulated time-domain pulse shapes of the corresponding platicon states. The fringes, marked with arrows, are related to the oscillating tails at the bottom of the dark pulse. d. Measured platicon’s repetition rate $f_{\text{rep}}$ (blue) and phase noise $S_{\phi}$ at 10 kHz Fourier offset frequency (red) of the 10.7-GHz microwave, with backward tuning. $f_{0}=10.68545$ GHz. Vertical dashed lines highlights that the local minimum of $S_{\phi}$ coincides with the local maximum of $f_{\text{rep}}$. Horizontal dashed lines highlights that this coincidence is associated with $\text{d}f_{\text{rep}}/\text{d}I=0$. e. Phase noise $S_{\phi}$ spectra of the local maximum and minimum points marked with green and black dots in d. Phase noise quenching up to 23.5 dB is observed. Inset shows the beatnote of the lowest $S_{\phi}$ (black data) with 10 Hz RBW.

### v2

@@
Shichang Li Shenzhen Institute for Quantum Science and Engineering, Southern University of Science and Technology, Shenzhen 518055, China International Quantum Academy, Shenzhen 518048, China    Kunpeng Yu International Quantum Academy, Shenzhen 518048, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China    Dmitry A. Chermoshentsev Russian Quantum Center, Moscow 143026, Russia    Wei Sun sunwei@iqasz.cn International Quantum Academy, Shenzhen 518048, China    Jinbao Long International Quantum Academy, Shenzhen 518048, China    Xiaoying Yan Shenzhen Institute for Quantum Science and Engineering, Southern University of Science and Technology, Shenzhen 518055, China International Quantum Academy, Shenzhen 518048, China    Chen Shen International Quantum Academy, Shenzhen 518048, China Qaleido Photonics, Shenzhen 518048, China    Artem E. Shitikov Russian Quantum Center, Moscow 143026, Russia    Nikita Yu. Dmitriev Russian Quantum Center, Moscow 143026, Russia    Igor A. Bilenko Russian Quantum Center, Moscow 143026, Russia Faculty of Physics, Lomonosov Moscow State University, Moscow 119991, Russia    Junqiu Liu liujq@iqasz.cn International Quantum Academy, Shenzhen 518048, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China
@@
Microcombs, formed in optical microresonators driven by continuous-wave lasers, are miniaturized optical frequency combs. Leveraging integrated photonics and laser self-injection locking (SIL), compact microcombs can be constructed via hybrid integration of a semiconductor laser with a chip-based microresonator. While the current linear SIL theory has successfully addressed the linear coupling between the laser cavity and the external microresonator, it fails to describe the complicated nonlinear processes, especially for dark-pulse microcomb formation. Here, we investigate – theoretically, numerically and experimentally – the Kerr-thermal dynamics of a semiconductor laser self-injection-locked to an integrated silicon nitride microresonator. We unveil intriguing yet universal dark-pulse formation and switching behaviour with discrete steps, and establish a theoretical model scrutinizing the synergy of laser-microresonator mutual coupling, Kerr nonlinearity and photo-thermal effect. Numerical simulation confirms the experimental result and identifies the origins. Exploiting this unique phenomenon, we showcase an application on low-noise photonic microwave generation with phase noise purified by 23.5 dB. Our study not only adds critical insight of pulse formation in laser-microresonator hybrid systems, but also enables all-passive, photonic-chip-based microwave oscillators with high spectral purity.
@@
Optical frequency combs (OFCs) [1, 2, 3, 4] have revolutionized timing, spectroscopy, precision measurement, and testing fundamental physics. Microcombs – harnessing enhanced Kerr nonlinearity in optical microresonators driven by continuous-wave (CW) pumps – have allowed miniaturized OFCs with small size, weight and power consumption [5, 6]. With the emergence and quick maturing of ultralow-loss photonic integrated circuits [7, 8, 9, 10], as well as heterogeneous and hybrid integration with semiconductor lasers [11, 12], today microcombs can be built entirely on-chip and manufactured in large volume with low cost, catalyzing wide deployment outside laboratories.
@@
Depending on the microresonator’s group velocity dispersion (GVD), there are two types of coherent microcombs. The bright dissipative soliton microcombs [13, 14, 15, 16, 17, 18] require anomalous GVD, while the dark-pulse (also termed “platicon”) microcombs require normal GVD [19, 20, 21, 22, 23, 24]. Compared to solitons, platicons exhibit remarkably higher CW-to-pulse power conversion efficiency [25, 26], thus are advantageous for coherent optical communication [27] and photonic microwave generation [28]. To initiate platicons, laser self-injection locking (SIL) [29, 30] offers the most robust and effective form, and permits seamless integration of the pump laser and the microresonator on a monolithic substrate [31, 32, 33]. However, while the current linear SIL theory and model [29] have addressed the linear coupling between the laser cavity and the external microresonator, it fails to manage the complicated nonlinear processes in platicon formation.
@@
Figure 1: Principle and schematic of SIL, Kerr nonlinearity and photo-thermal effect in the laser-microresonator coupled system. a. Frequency- and time-domain picture describing the synergy of SIL, photo-thermal effect and Kerr nonlinearity, which yields narrowing of the pump laser’s linewidth, red-shift of the microresonator’s resonance grid, and formation of a platicon microcomb. Res. mode: resonance mode. b. Experimental setup. A DFB laser chip edge-coupled to a Si3N4 microresonator chip. A printed circuit board (PCB) provides stable current to the laser and stabilizes the laser temperature. Output light from the Si3N4 microresonator is collected by a lensed fiber. $E_{\mathrm{L}}$: light amplitude in the laser cavity. $E^{+}$/$E^{-}$: light amplitude in the clockwise/counter-clockwise direction in the microresonator. $E_{\mathrm{out}}$: light amplitude at the microresonator output. $\varphi$: feedback phase of $E^{-}$.
@@
Here, we investigate the Kerr-thermal dynamics of platicon formation using SIL of a semiconductor laser to an integrated silicon nitride (Si3N4) microresonator. We unveil an intriguing platicon switching behaviour with discrete steps, allowing operation of platicons in quiet states immune to laser noise. Figure 1a illustrates the principle of our study. When a laser-cavity mode is tuned into a resonance mode of the Si3N4 microresonator and light is coupled into the microresonator, SIL occurs that locks the laser frequency to the resonance [29, 30] and suppresses the laser linewidth. The linewidth suppression ratio is proportional to $Q_{\mathrm{r}}^{2}/Q_{\mathrm{d}}^{2}$, where $Q_{\mathrm{r/d}}$ is the quality factor of the microresonator/laser cavity. Meanwhile, photo-thermal effect [34, 35, 36, 37] induces a global frequency shift of the microresonator’s resonance grid. With sufficient intracavity power, Kerr nonlinearity induces four-wave mixing (FWM) [38] that translates photons to other resonances. In microresonators of normal GVD, the synergy of SIL and Kerr effect ultimately yield platicon formation.
@@
Experimental result. We use a commercial DFB laser chip edge-coupled to a Si3N4 microresonator chip [39, 28], as shown in Fig. 1b. The DFB laser operates at 1549 nm. It has 1 nm wavelength tuning range and 159 mW output CW power with driving current up to 500 mA. The Si3N4 microresonator has 10.7 GHz free spectral range (FSR) and normal GVD ($D_{2}<0$). The microresonator is critically coupled [40], with intrinsic quality factor $Q_{0}=23\times 10^{6}$. Actual photographs and characterization data of the DFB lasers and the Si3N4 microresonators are found in Supplementary Information Note 1.
@@
Light $E_{\mathrm{L}}$ in the laser cavity is emitted and coupled into the microresonator’s clockwise direction $E^{+}$. Rayleigh scattering in the microresonator reflects some light to the counter-clockwise direction $E^{-}$ and to the laser cavity. To facilitate SIL, we optimize the edge coupling and the gap distance (thus the feedback phase $\varphi$) between the two chips. In the SIL regime, we continuously tune the laser current, and monitor the output optical power and frequency from the Si3N4 chip. The output light beats against a reference CW laser, and the beat frequency is recorded. The forward (backward) tuning corresponds to increasing (decreasing) laser current, which decreases (increases) laser frequency.
@@
Figure 2(a, b) shows the experimental data. When SIL occurs, the output power experiences a sudden drop, and the output frequency is shifted. In the SIL regime, we observe synchronous dynamics of the output power and the frequency, due to the photo-thermal effect [34, 35, 36, 37]. For example, the decreasing output power corresponds to decreasing transmitted power and increasing intracavity power. The latter causes an increasing resonance red-shift (towards longer wavelength) via the photo-thermal effect[34]. As the laser frequency is locked to the resonance, the red-shifted resonance drags the laser frequency, leading to decreasing output frequency.
@@
Besides, it is apparent that the output power and frequency exhibit discrete step features. This is contrary to the conventional linear SIL model [29] showing continuous, nearly anchored tuning curves of the frequency. Here, the appearance of these steps is attributed to the Kerr nonlinearity in the microresonator with sufficient intracavity power. Note that similar discontinuous curves have been observed and characterized in self-injection-locked solitons [31], while here we observe and characterize these curves for the first time for platicons. Figure 2c shows the zoom-in profile of the gray-shaded zone in Fig. 2b. Steps in backward tuning with 320.4 to 321.3 mA laser current are marked with 1 to 4. Typical optical spectra within each step are shown in Fig. 3a, evidencing formation and switching of different platicon states. Distinct fringes are marked with arrows on the spectral envelopes. Numerical simulation of time-domain pulse shapes in Fig. 3a (discussed later) confirms that all these platicon states are “single platicons” comprising only one dark pulse in the microresonator.
@@
Moreover, identical Kerr-thermal platicon dynamics has also been observed in other two independent SIL setups, where different DFB lasers and a 21.3-GHz-FSR Si3N4 microresonator are used. Details are found in Supplementary Information Note 1 and 2. These parallel experiments suggest that, our observation of the Kerr-thermal dynamics and discrete steps is universal and independent of the particular devices.
@@
Figure 2: Experimental result in comparison with numerical simulation. a. Experimentally measured output laser power with forward (green) or backward (yellow) tuning. b. Experimentally measured beat frequency between the output laser and a frequency-fixed reference laser, with forward (blue) or backward (red) tuning. c. Zoom-in profiles of the gray-shaded region in b with forward (blue) or backward (red) tuning. Discrete steps with backward tuning are numbered with 1 to 4. d, e. Numerical simulation results corresponding to experimental data in a, b. Horizontal axes are the frequency detuning (in the unit of $\kappa/2$) between the free-running laser frequency and the cold resonance frequency. Effective (Eff.) detuning is the frequency detuning (in the unit of $\kappa/2$) between the laser frequency at the microresonator output and the cold resonance frequency. f. Zoom-in profile of the gray-shaded zoom in d, e. Gray curves outline the full detuning range of platicon steps.
@@
Theory and numerical simulation. To better understand our experimental result, we establish a comprehensive theoretical model synergizing SIL, Kerr nonlinearity, and photo-thermal effect. The equations describing our laser-microresonator nonlinear system are written as
$\partial_{\tau}N$ $=\gamma_{\mathrm{e}}\left[\gamma-F(N-1)-(N-1+\gamma)|a_{\mathrm{L}}|^{2}\right]$ (1)
@@
$\partial_{\tau}a_{\mathrm{L}}$ $=\frac{1}{2}\gamma_{\mathrm{p}}\left(1+i\alpha_{\mathrm{g}}\right)(N-1)a_{\mathrm{L}}+i\mathrm{e}^{i(\alpha_{\mathrm{L}}\tau+\varphi)}\frac{\kappa_{\mathrm{c}}}{f_{\mathrm{e}}}b_{0}$ (2)
@@
$\partial_{\tau}a_{\mu}$ $=-\left(1+i\sum_{n=2}^{\infty}\frac{d_{n}}{n!}\mu^{n}-iT\right)a_{\mu}+i\left(\mathcal{F}[|a|^{2}a]_{\mu}+2a_{\mu}\sum_{\mu^{\prime}}\left|b_{\mu^{\prime}}\right|^{2}\right)+i\tilde{\beta}b_{\mu}+i\mathrm{e}^{i(-\alpha_{\mathrm{L}}\tau+\varphi)}f_{\mathrm{e}}a_{\mathrm{L}}\delta_{0\mu}$ (3)
@@
$\partial_{\tau}T$ $=\left[k_{T}\sum_{\mu}\left(\left|a_{\mu}\right|^{2}+\left|b_{\mu}\right|^{2}\right)-T\right]/t_{\mathrm{h}}.$ (5)
@@
Derivation of Eq. (1–5) is found in Supplementary Information Note 3. In short, Eq. (1) describes the time evolution of the electron carrier density $N$ normalized to the threshold carrier density inside the laser cavity, with $\tau=\kappa t/2$, $\gamma_{\mathrm{e}}=[2I/(\kappa eV_{\mathrm{L}}N_{0})-2\kappa_{N}/\kappa]/\gamma$, $\gamma=G_{0}/(G_{\mathrm{n}}N_{0})$, $F=2\kappa_{N}/\kappa/\gamma_{\mathrm{e}}$. Here, $\kappa/2\pi$ is the loaded linewidth of the Si3N4 microresonator, $t$ is the real time, $I$ is the driving current on the laser diode, $e$ is the electron charge, $V_{\mathrm{L}}$ is the effective volume of the laser diode’s active area, $\kappa_{N}$ is the dissipation rate of carrier density, $N_{0}$ is the electron carrier density of the free-running laser with a fixed driving current, $G_{0}$ is the gain of the light intensity per second at $N_{0}$ in the laser cavity, $G_{\mathrm{n}}$ is the gain coefficient per second, $a_{\mathrm{L}}$ represents the light amplitude in the laser cavity.
@@
Equation (2) describes the laser dynamics in the laser cavity, with $\gamma_{\mathrm{p}}=2G_{\mathrm{n}}N_{0}/\kappa$, $\alpha_{\mathrm{L}}=2\delta\omega_{\mathrm{L}}/\kappa$, $\varphi=\omega_{0}t_{\mathrm{s}}$, $\kappa_{\mathrm{c}}=4\kappa_{\mathrm{ex}}T_{\mathrm{L}}\sqrt{T_{\mathrm{cL}}T_{\mathrm{cR}}}/(\kappa^{2}\tau_{\mathrm{L}})$, $f_{\mathrm{e}}=2\sqrt{\gamma_{\mathrm{e}}\kappa_{\mathrm{ex}}T_{\mathrm{L}}T_{\mathrm{cR}}gV_{\mathrm{L}}}/\sqrt{\kappa^{2}G_{\mathrm{n}}\tau_{\mathrm{L}}}$. Here, $\delta\omega_{\mathrm{L}}$ represents the frequency difference between the cold-cavity resonance mode $\omega_{\mathrm{L}}$ and the microresonator resonance mode $\omega_{0}$, $t_{\mathrm{s}}$ is the time delay between the laser cavity and the microresonator, $\kappa_{\mathrm{ex}}$ is the external coupling rate of the microresonator, $T_{\mathrm{L}}$ is the light transmission efficiency at the laser cavity’s output surface, $T_{\mathrm{cR}}$ is the light coupling efficiency from the laser cavity to the microresonator, $T_{\mathrm{cL}}$ is the light coupling efficiency from outside into the laser cavity, $\tau_{\mathrm{L}}$ is the light round-trip time in the laser cavity, $g$ is the nonlinear coefficient of the Si3N4 microresonator. $\alpha_{\mathrm{g}}$ is the linewidth enhancement factor and $b_{0}$ represents the back-scattered light in the microresonator at the $0^{\mathrm{th}}$ resonance mode.
@@
Equations (3, 4) describe the dynamics of the forward-propagating light $a_{\mu}$ and the back-scattered light $b_{\mu}$ at the $\mu^{\mathrm{th}}$ resonance mode, with $d_{n}=2D_{n}/\kappa$, and $\tilde{\beta}=2\beta/\kappa$. Here, $D_{n}$ is the $n^{\mathrm{th}}$-order dispersion of the microresonator, $\beta$ is the back-scattering ratio between the back-scattered light and the forward-propagating light, and $\mathcal{F}[|a|^{2}a]$ ($\mathcal{F}[|b|^{2}b])$ represents the Kerr interaction.
@@
Equation (5) describes the evolution of the normalized frequency shift $T$ induced by the temperature in the microresonator, with $k_{T}=K_{T}/g$, and $t_{\mathrm{h}}=\kappa\tau_{\mathrm{h}}/2$. Here, $K_{T}$ is the thermal-induced resonance shift coefficient, and $\tau_{\mathrm{h}}$ is the thermal relaxation time.
@@
Figure 3: Optical spectra of different platicon states and the resulted noise-quenching effect. a. Experimentally measured (blue lines) and simulated (red curves) optical spectra of the platicon states labeled in Fig. 2c, f. They agree not only on the spectral envelopes but also on the number of fringes (marked with arrows). Insets: simulated time-domain pulse shapes of the corresponding platicon states. The fringes, marked with arrows, are related to the oscillating tails at the bottom of the dark pulse. b. Measured platicon’s repetition rate $f_{\mathrm{rep}}$ (blue) and phase noise $S_{\phi}$ at 10 kHz Fourier offset frequency (red) of the 10.7-GHz microwave, with backward tuning. $f_{0}=10.68545$ GHz. Vertical dashed lines highlight that the local minimum of $S_{\phi}$ coincides with the local maximum of $f_{\mathrm{rep}}$. Horizontal dashed lines highlight that this coincidence is associated with $\mathrm{d}f_{\mathrm{rep}}/\mathrm{d}I=0$. c. Phase noise $S_{\phi}$ spectra of the local maximum and minimum points marked with green and black dots in b. Phase noise quenching up to 23.5 dB is observed. Inset shows the beat note of the lowest $S_{\phi}$ (black data) with 10 Hz RBW.